OpenStudy (anonymous):

(-4x^2y)^3 I really need help, I still don't understand.

OpenStudy (anonymous):

\[ (-4x^{2y})^{3} is this what you mean?\]

OpenStudy (anonymous):

\[(ab^{c})^{d} = a^{d} b ^{cd}\]

OpenStudy (anonymous):

so it will be \[-4^{3}x ^{6y}\]

OpenStudy (anonymous):

= -64x^6y

OpenStudy (anonymous):

remember that y has an implicit exponent of 1, and 1*3 = 3.

OpenStudy (anonymous):

no i meant (-4x to the second power than y)^3 the y is not part of the exponent

OpenStudy (anonymous):

oh, so it's \[(-4x ^{2}y)^{?}\]

OpenStudy (anonymous):

the concept is the same, -4^3 x^6 y^3

OpenStudy (anonymous):

or -64x^6 y^3

OpenStudy (anonymous):

I think the question is :\[(-4x^2y)^3\] right?

OpenStudy (anonymous):

yes thats right, i just dont understand how to solve it

OpenStudy (anonymous):

As I said, when raising a product to a power, you multiply the exponents of the factors. So the 3 would be multiplied to the power of the -4 factor, and the power of the x factor, and the power of the y factor.

OpenStudy (anonymous):

alright, first of all you've got to mulitply the power 3 with the given between the brackets and you'll get ^_^: \[=(-4)^3 x^6 y^3 = 48x^6y^3\] is it clearer now ? :)

OpenStudy (anonymous):

\[(-4)^3 \ne 48 \] \[(-4)^3 = -64 \]

OpenStudy (anonymous):

lol! sry >_< silly mistake , thank you polpak :)

OpenStudy (anonymous):

do you understand it now courtney ^_^?

OpenStudy (anonymous):

just substitute the power to each :)

OpenStudy (anonymous):

just pass the power out to each*

OpenStudy (anonymous):

yes, kind of but there is another problem now that is in the same format but is a fraction and has that equal sign with a / through it. grr i hate math lol

OpenStudy (anonymous):

don't say that, look at the problem , write it down on paper and try to use tha same concept like the one here :) The problem never gets harder, it just gets more detailed ^_^

OpenStudy (anonymous):

nothing changes, just apply the same rule :)

OpenStudy (anonymous):

The rule applies equally to division. The only part that gets messy is when you have a sum that you're raising to a power.

OpenStudy (anonymous):

im trying to stay positive, thank you so much for your help you two are life savers! im writing all the work you just sent as my work shown but what part is the part that i put in the space for the answer to the problem?

OpenStudy (anonymous):

Try to see if you can figure it out from what we've said. We'll check your answer.

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

ok im working it out now than ill post it on here, thank you again.

OpenStudy (anonymous):

\[(((3p ^{3}v ^{4} \ s ^{4}))^{2}\]

OpenStudy (anonymous):

ok this is a fraction the s^4 is on the bottom and it also says s with the equal sign / 0

OpenStudy (anonymous):

i got 9p^6y^8 over s^8

OpenStudy (anonymous):

So.. \[(3p^3v^4/s^4)^2, s \ne 0\]

OpenStudy (anonymous):

Then yes, that is the correct answer. Good job!

OpenStudy (anonymous):

yayyy thank you, even my professor cant teach me math and over the computer you got right into my head lol

OpenStudy (anonymous):

It may be somewhat helpful to expand the exponents for a while to see what is happening. \[(3p^3v^4/s^4)^2 = [(3pppvvvv)/(ssss)]^2 \] \[= [(3pppvvvv)(3pppvvvv)]/[(ssss)(ssss)]\] \[= [(3*3)ppppppvvvvvvvv]/[ssssssss]\] \[= 9p^6v^8/s^8\]

OpenStudy (anonymous):

that acutally helps alot, can i ask you on more question?

OpenStudy (anonymous):

Of course!

OpenStudy (anonymous):

ok so i know how to do this in a normal way but they want it in exponential form and im confused. the first is (5^3)^3 and the next is in fraction form which is even worse lol (r/s)^6 (s\[\neq\]0)

OpenStudy (anonymous):

So try to expand it out. \[(5^3)^3 = (5^3)(5^3)(5^3)\] Now how would you expand that further?

OpenStudy (anonymous):

five times five times five than add the exponents?

OpenStudy (anonymous):

\[5^3 = 5*5*5\] \[\rightarrow 5^3*5^3*5^3 = (5*5*5)*(5*5*5)*(5*5*5)\] \[=5^?\]

OpenStudy (anonymous):

5^9?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

awesome now to the fraction, can you help me with that one?

OpenStudy (anonymous):

Now we can see that the faster way to arrive at the same result would be to simply : \[(5^3)^3 = 5^{3*3} = 5^9\]

OpenStudy (anonymous):

okay so you multiply the exponents too?

OpenStudy (anonymous):

Well, remember when we expanded it out, we said that we had three groups with three 5's each. So how many 5's do we have multiplied together? 3*3 right?

OpenStudy (anonymous):

right, im trying to keep up im just so lost

OpenStudy (anonymous):

So if you're not comfortable with multiplying the exponents, just try expanding the next one out for yourself. \[a^6 = a*a*a*a*a*a\] \[ \rightarrow (r/s)^6 =\ ?\]

OpenStudy (anonymous):

We'll just go one step at a time.

OpenStudy (anonymous):

would i seperate them and do r*r*r*r*r*r and s*s*s*s*s*s?

OpenStudy (anonymous):

No, just multiply the whole thing.

OpenStudy (anonymous):

so r/s*r/s*r/s*r/s*r/s*r/s

OpenStudy (anonymous):

Yes, but keep the parens so it's more readable. \[(r/s)*(r/s)*(r/s)*(r/s)*(r/s)*(r/s)\] And when we multiply two fractions, we do what?

OpenStudy (anonymous):

top times top bottom times bottom?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so its (r^6)/(s^6)

OpenStudy (anonymous):

Yes. And here again we can arrive at this answer in a more direct way once we understand the process \[ (r/s)^6 = (r^1/s^1)^6 = r^{1*6}/s^{1*6} = r^6/s^6\]

OpenStudy (anonymous):

that would be the answer correct since we do not have the values for r and s? what about the =/ thing?

OpenStudy (anonymous):

They have to make the restriction \[s \ne 0\] because if s is 0 then you're dividing by 0 which is not defined. That answer is correct regardless of what r and s is. If we know more about them we might be able to simplify further (if s was 2 and r was 4 for example). But even in that case our answer is correct. It just might not be in the form they're looking for.

OpenStudy (anonymous):

yayyy you are so amazing and kind hearted. that you so so much!

OpenStudy (anonymous):

You're very welcome.